Twelve different video games showing substance use were observed and the duration times of game play​ (in seconds) are listed below. The design of the study justifies the assumption that the sample can be treated as a simple random sample. Use the data to construct a 90​% confidence interval estimate of mu​, the mean duration of game play.
4040
3881
3858
4031
4328
4820
4658
4042
4998
4821
4344
4319

Click here to view a t distribution table.LOADING...
Click here to view page 1 of the standard normal distribution table.LOADING...
Click here to view page 2 of the standard normal distribution table.LOADING...
What is the confidence interval estimate of the population mean mu​?
nothingless thanmuless than
nothing
​(Round to one decimal place as​ needed.)
Answer #1

sample mean, xbar = 4345
sample standard deviation, s = 394.6736
sample size, n = 12
degrees of freedom, df = n - 1 = 11

Given CI level is 90%, hence α = 1 - 0.9 = 0.1
α/2 = 0.1/2 = 0.05, tc = t(α/2, df) = 1.796


CI = (xbar - tc * s/sqrt(n) , xbar + tc * s/sqrt(n))
CI = (4345 - 1.796 * 394.6736/sqrt(12) , 4345 + 1.796 * 394.6736/sqrt(12))
CI = (4140.4 , 4549.6)